# 111preprint

### v1

@@ 4. Estimated Human Behavioral Model
To estimate the human behavioral and trust dynamics model parameters, a human subject experiment is conducted333 The human behavioral experiments were approved under Michigan State University Institutional Review Board Study ID 9452.. A total of 39 participants were recruited, comprising 17 males and 22 females, all of whom were freshman or sophomore college students and the majority were not in the engineering program. The experiment lasted on average for 1 hour and participants are compensated with an amount of \$$\$15.00$ for their participation. Participants completed 71 trials in total, with 41 trials in low complexity and the remaining 30 trials in high complexity.
When the robot attempts to collect an object autonomously, it has a success probability of $p^{\textup{suc}}_{H}=0.75$ in high complexity and $p^{\textup{suc}}_{L}=0.97$ in low complexity environment. These success rates are chosen given the expectation that low-complexity trials are deemed easier as compared with high-complexity trials. In each trial, the robot requests human assistance with a probability of $0.10$ in a low-complexity environment and with a probability of $0.33$ in a high-complexity environment.
@@
Before starting the experiment, participants receive a description outlining the task, the choices of robot actions, and the supervisor’s role. Following this, participants watch a video of the robot operating and collecting objects. They are informed that the robot is able to recognize the complexity with high probability and adjust itself to collect the cans safely, and it may occasionally ask for supervisor assistance. They then proceed to a familiarization phase in a practice environment where they are faced with all possible combinations of complexity and robot actions. They are asked to occasionally interrupt and intervene to get familiarized with how to interrupt and control the manipulator using the joystick.
@@
The collected data in each trial are human action $o$, robot action $a$, complexity $C$, and experience $E$. The data collected from all participants are combined and used to estimate the parameters of the human behavioral and trust dynamics model. Each time a shelf is cleared and objects are collected, human participants are asked to fill out a survey related to their trust. It should be noted that the survey responses are not used for learning the parameters of the human behavior model. The survey responses collected are discussed in Section 4.4. To learn the parameters of the IOHMM-based human behavior model, an extended version of the Baum-Welch algorithm is used(536317, ). The estimated parameters of the model are presented and discussed in the following section. For brevity and simplification of notation, the events are represented by their outcomes; for example, $E_{t}=E^{+}$ is shortened to $E^{+}_{t}$.
Since the data was collected through in-person experiments, we were able to observe the behaviors of the participants and monitor them. Data collected from 6 participants are omitted due to behaviors observed during the experiment such as lack of attention, constantly checking their phone, and unable to follow the basic instructions of the experiment. This leaves 33 data sequences for model parameter estimation.
@@
(c) [$E^{-}$, $C^{L}$, $a^{-}$]
Figure 5. Estimated trust state transition probabilities $\mathbb{P}(T_{t+1}|T_{t},E_{t+1},C_{t},a_{t})$ for low environment complexity $C_{t}=C^{L}$
[…]
(a) [$E^{+}$, $C^{H}$, $a^{+}$]
[…]
(b) [$E^{-}$, $C^{H}$, $a^{+}$]
[…]
(c) [$E^{+}$, $C^{H}$, $a^{-}$]
Figure 6. Estimated trust state transition probabilities $\mathbb{P}(T_{t+1}|T_{t},E_{t+1},C_{t},a_{t})$ for high environment complexity $C_{t}=C^{H}$
4.1. Initial Trust and Observation Probabilities.
The estimated initial probability of trust being high is $\mathbb{P}(T^{H})=0.82$. This estimate suggests that at the beginning of each task, humans are expected to have a relatively high trust towards the robot.
The estimated observation probabilities $\mathbb{P}(o^{+}_{t}|T_{t},C^{L}_{t},a^{+}_{t})$ is shown in Fig. 4(b). In low complexity, when the robot does not ask for human assistance and attempts to collect autonomously, the human relies with probability 1 and 0.97 when their trust is high and low respectively. Irrespective of the human trust level, there is a higher probability of reliance in low complexity scenarios due to the higher success rate, with the supervisor perceiving lower risk in such situations.
The estimated observation probabilities $\mathbb{P}(o^{+}_{t}|T_{t},C^{H}_{t},a^{+}_{t})$ is shown in Fig. 4(c). In high complexity, when the robot does not ask for human assistance and attempts to collect autonomously, the human relies with probability 0.94 and 0.43 when their trust is high and low respectively. As compared with the parameters obtained in low complexity, the probability of reliance in high complexity is lower in general. The high probability of reliance, when their trust is high, is indicative of humans perceiving the robot to be efficient and reliable; while the high probability of interruption, when their trust is low, is indicative of humans perceiving a higher risk and possible failure in high complexity.
4.2. State Transition Matrices.
The estimates obtained for trust state transition in the low complexity environment are shown in Fig. 5. The estimates for $\mathbb{P}(T_{t+1}|T_{t},E^{+}_{t+1},C^{L}_{t},a^{+}_{t})$ can be seen in Fig. 5(a). If the current trust is high, it remains high with a probability of 1; conversely, if the current trust is low, it remains low with a probability of 1. Success in low complexity does not help in repairing trust, showing that success in an easy task has very little to no effect on restoring trust.
The estimates for trust transition when there’s a failure in low complexity, $\mathbb{P}(T_{t+1}|T_{t},E^{-}_{t+1},C^{L}_{t},a^{+}_{t})$, is shown in Fig. 5(b). In this scenario, if the trust is currently high, it transitions to low with a probability of 0.71; while if the current trust is low, it remains low with a probability equal to 1. This shows that failure negatively impacts trust, which is to be expected.
We will now discuss the state transition when operating in high complexity. Fig. 6 shows the estimates of state transition probabilities for high environment complexity. Fig. 6(a) shows the model estimates for a success in autonomous collection $\mathbb{P}(T_{t+1}|T_{t},E^{+}_{t+1},C^{H}_{t},a^{+}_{t})$. It can be seen that if the current trust is high, the trust level stays high with probability 1, and the trust transitions from low to high with probability 0.64. Thus, success in high-complexity tasks can repair trust.
The estimate of $\mathbb{P}(T_{t+1}|T_{t},E^{+}_{t+1},C^{H}_{t},a^{+}_{t})$, the case when autonomous collection fails, is shown in Fig. 6(b). Human trust transitions from high to low with probability equal to 0.33 and remains in low with probability 0.88. Thus, as expected, a failure in high complexity has a negative effect on trust. With the reduced autonomous success rate in high complexity, these tasks are high risk and high reward. In summary, a failure negatively impacts the trust level, and success in more complex tasks $C^{H}$ has a higher chance of repairing trust as compared to success in an easier task $C^{L}$.
@@ 4.3. Model Parameter Confidence
We now investigate uncertainty in the estimated parameters. Since the Baum-Welch algorithm provides a maximum a posteriori estimate, we rely on Laplace approximation (pml2Book, ) to estimate parameter uncertainty. To this end, we estimate the Hessian matrix $H$ of the log-likelihood $L$ and compute the uncertainty in parameter estimates as $e=\sqrt{\operatorname{diag}(-H)}$, where $\operatorname{diag}$ represents the diagonal elements of the matrix $-H$. The uncertainties in the estimated IOHMM human trust model parameters are shown in Table 1.
@@
For parameters related to $C^{L}$, the standard errors are relatively low and accurate up to the first significant digit of the estimate, with no more than ±$\pm 0.9$ error in the parameter. Specifically, this refers to the human action probability and $\mathbb{P}(T^{H}|T,E^{+},a^{+})$ parameters.However, the errors for other parameters are significantly higher, with the highest having an error of ±$\pm 0.56$. These high-uncertainty parameters are associated with scenarios involving failures and when the robot asks for assistance and can be explained by the low frequency of these scenarios during the experiment.
The parameters conditioned on the state being in high trust $T^{H}$ have low standard errors, with the highest standard error being ±$\pm 0.07$. This suggests that these parameter estimates are accurate up to the first significant digit. In contrast, the standard errors for parameters conditioned on the state being in low trust $T^{L}$ are relatively high, ranging from ±$\pm 0.15$ to ±$\pm 0.3$. This high uncertainty can possibly be explained by the model’s difficulty in accurately capturing/encountering the low trust state, as compared to the high trust state.
@@ 4.4. Trust as a Hidden State
We now investigate if it is justified to refer to the hidden state in our model as “human trust”. During the experiment, participants completed a survey (lee1992trust, ) in which they responded, on an 11-point scale, to the following question “To what extent do you trust the automation in this scenario?”, shown in Appendix 8.2. This survey is collected each time the robot finishes collecting objects from a shelf, before moving to another shelf. Using the model parameters, we calculate the probability of being in a particular state at each time step via the belief update. Given the assumption that trust evolution follows a Markov process, the probability of trust being high at the time indices when a shelf is cleared can be considered as an aggregated representation of the trust state based on previous interactions. We then compare the probability of trust being high at the end of each shelf-clearing event to the corresponding survey responses
The comparison of the probability of trust being high computed using the estimated model, and survey responses is shown in Fig. 8. It can be observed that the median probability of trust being high increases with trust scores collected through the survey. While the mid-level trust scores have higher variance, the higher trust estimates have lower variance in their trust response. This can signify that humans are unsure about their trust score response when their trust is low as compared to when their trust is high. Lower trust levels (0,1, and 2) did not receive any rating from participants. This can indicate that participants are more certain when their trust is high but uncertain about their trust level when it is low. The median of the probabilities of trust being high is used to fit a curve following a logistic function resulting in
@@
where the function $P^{T}(r^{T})$ represents the probability of high trust as a function of the human trust response $r^{T}$. The curve is fitted with $95\%$ confidence bounds using MATLAB’s fit function. The plot of $P^{T}(r^{T})$ with the coefficients as compared with the median values of high trust estimate per survey response can be seen in Fig. 8. The results suggest that the hidden state in our model is isomorphic to self-reported trust scores, and the probability of trust being high is a sigmoidal function of the self-reported trust values.
@@ 4.5. Trust and Human Reliance
It can be observed that the median survey responses can be approximated as a linear function of the trust survey response. The estimated linear function which maps trust response to reliance response can be approximated by the function
@@ 5.1. Computed Assistance-seeking Policy
The baseline policy that will be used is a trust-agnostic policy which does not take the trust state into account. For the trust-agnostic policy, the only attribute of a trial is the environmental complexity. We used the data from the experiment in Section 4 to estimate the probability $\mathbb{P}(o_{t}=o^{-}|C_{t},a^{-})$, which is calculated by taking the empirical estimate in the dataset as
@@
$\mathbb{E}[R|a^{-},C]$ $=\mathcal{R}^{a^{+}}_{o^{+},E^{+}}\mathbb{P}(o_{t}=o^{+}|C_{t},a^{+})p^{\textup{suc}}_{C}$
$+R^{a^{+}}_{o^{+},E^{-}}\mathbb{P}(o_{t}=o^{+}|C_{t},a^{+})(1-p^{\textup{suc}}_{C})$
$+R^{a^{+}}_{o^{-}}\mathbb{P}(o_{t}=o^{-}|C_{t},a^{+}),$
@@
Since the complexity of the environment follows an i.i.d random variable, for trust-agnostic policy, it suffices to simply calculate the expected reward per action for a given complexity to obtain the optimal action. With the data collected and our experiment design and parameters, the estimated rewards for attempting to collect autonomously are $\mathbb{E}[R|a^{+},C^{L}]=2.79$ and $\mathbb{E}[R|a^{+},C^{H}]=1.02$. Comparing the expected rewards for each action in the respective complexities, these calculations suggested that the optimal trust-agnostic policy is to never seek assistance irrespective of the environmental complexity, $a(t)=a^{+}$ for each $t$.
@@ 5.2. Evaluation of the Proposed Policy
To evaluate the performance of the calculated optimal trust-aware assistance-seeking policy, a second set of experiments with human participants is conducted. A total of 20 participants were recruited and participated in the experiment, each completing two blocks of experiments in which the robot followed trust-aware and trust-agnostic policies, respectively. The participants recruited are different from those recruited in the first set of experiments. The order of the blocks was randomly assigned for each participant, with 10 participants starting with the trust-aware policy and the other 10 starting with the trust-agnostic policy. In each block, participants performed 20 low-complexity and 20 high-complexity trials. Due to the same reason as for the first experiment, data from 2 participants was omitted, leaving us with 18 data sequences. The cumulative reward for all participants under both policies is illustrated in Fig. 11. The usage of trust-agnostic policy resulted in a median score of 69, while the trust-aware policy resulted in a median score of 84. The trust-aware policy collected more objects autonomously and prevented more human interruptions as compared with the trust-agnostic policy. A t-test with $5\%$ significance level shows that trust-aware policy outperforms trust-agnostic policy with a p-value of $p=0.0007$, and is calculated using MATLAB’s ttest2 function.
Figure 11. Cumulative reward statistics for the trust-agnostic and trust-aware policies. The trust-aware policy outperformed the trust-agnostic policy with a p-value of: 0.0007
5.3. Survey Based Evaluation of the Trust-Aware Policy
@@
In total, participants completed this questionnaire three times, the first before the experiment, and the remaining two after completion with each policy. We dub these survey responses as “pre-experiment”, “post-trust-agnostic”, and “post-trust-aware” responses. The comparison of the survey scores is shown in Fig. 12. The responses to the 14 survey questions are aggregated, resulting in a trust score. This process involves first reverse coding (taking the negative value) the questions that describe negative qualities. The responses are then summed and divided by the total number of questions. More detailed instructions can be found in (schaefer2016measuring, ). The average scores on the collected trust survey are 44, 44.9, and 50.7 for pre-experiment, trust-agnostic, and trust-aware, respectively. Table 2 presents the calculated p-values associated with survey data collected under different experiment conditions. When comparing surveys collected after the experiment, the trust-aware policy (mean score of 50.7) yields a higher trust score relative to the trust-agnostic policy (mean score of 44.9) with a p-value of 0.0342, which is statistically significant. When comparing surveys collected before the experiment (mean score of 44) and post-trust-agnostic policy (mean score of 44.9), the p-value calculated is equal to 0.7722, which suggests that trust-agnostic policy has no statistically significant effect on trust. A comparison of the surveys collected before the experiment (mean score of 44) and after the experiment with the trust-aware policy (mean score of 50.7) suggests that the trust-aware policy improves human trust and the increase is somewhat statistically significant with a p-value of 0.0554.
Table 2. t-test results showing the p-values, comparing the collected trust surveys before the experiment and after the experiment with the trust-aware and trust-agnostic policy.
Surveys p-value
pre-experiment and post-trust agnostic 0.7722
pre-experiment and post-trust aware 0.0554
post-trust aware and post-trust agnostic 0.0342
Figure 12. Comparison of the collected trust surveys before the experiment and after the experiment with the trust-aware and trust-agnostic policy.

### v2

@@ 4. Estimated Human Behavioral Model
To estimate the human behavioral and trust dynamics model parameters, a human subject experiment is conducted [^2] The human behavioral experiments were approved under Michigan State University Institutional Review Board Study ID 9452.. A total of 39 participants were recruited, comprising 17 males and 22 females, all of whom were freshman or sophomore college students and the majority were not in the engineering program. To elicit human trust, we have informed the participants that they can receive compensation of up to $\mathdollar 15$, depending on their performance in the experiment. The experiment lasted on average for 1 hour and participants are compensated with an amount of $\mathdollar 15.00$ for their participation, regardless of their performance. While all participants ultimately received the full $\mathdollar 15$, this was not disclosed until after the study. This design element was intended to introduce a perceived risk of loss and thereby establish conditions under which trust could highly influence participants’ decisions. Participants completed 71 trials in total, with 41 trials in low complexity and the remaining 30 trials in high complexity. When the robot attempts to collect an object autonomously, it has a success probability of $p^{\textup{suc}}_{H}=0.75$ in high complexity and $p^{\textup{suc}}_{L}=0.97$ in low complexity environment. These success rates are chosen given the expectation that low-complexity trials are deemed easier as compared with high-complexity trials. In each trial, the robot requests human assistance with a probability of $0.10$ in a low-complexity environment and with a probability of $0.33$ in a high-complexity environment.
@@
Before starting the experiment, participants received a description outlining the task, the choices of robot actions, and the supervisor’s role. Following this, participants watch a video of the robot operating and collecting objects. They are informed that the robot is able to recognize the complexity with high probability and adjust itself to collect the objects safely, and it may occasionally ask for supervisor assistance. They then proceed to a familiarization phase in a practice environment where they are faced with all possible combinations of complexity and robot actions. They are asked to occasionally interrupt and intervene to get familiarized with controlling the manipulator using the joystick.
@@
The collected data in each trial are human action $o$, robot action $a$, complexity $C$, and experience $E$. The data collected from all participants are combined and used to estimate the parameters of the human behavioral and trust dynamics model. Each time a shelf is cleared and objects are collected, human participants are asked to fill out a survey related to their trust. It should be noted that the survey responses are not used for learning the parameters of the human behavior model. The survey responses collected are discussed in Section 4.4. To learn the parameters of the IOHMM-based human behavior model, an extended version of the Baum-Welch algorithm is used (536317, ). [^3] Our code is available at https://github.com/donghaem14/Learning-algorithm-for-Input-Output-HMM-with-state-and-observation-inputs. The estimated parameters of the model are presented and discussed in the following section. For brevity and simplification of notation, the events are represented by their outcomes; for example, $E_{t}=E^{+}$ is shortened to $E^{+}_{t}$. Since the data was collected through in-person experiments, we were able to observe the behaviors of the participants and monitor them. Data collected from 6 participants are omitted due to behaviors observed during the experiment such as lack of attention, constantly checking their phone, and inability to follow the basic instructions of the experiment. This leaves 33 data sequences for model parameter estimation.
(a) $a=a^{-}$, $\forall C$
(b) $a=a^{+}$, $C=C^{L}$
(c) $a=a^{+}$, $C=C^{H}$
Figure 4. Estimated observation probabilities $\mathbb{P}(o_{t}|T,C,a)$:
(a) $E^{+}$, $C^{L}$, $a^{+}$]
(b) $E^{-}$, $C^{L}$, $a^{+}$]
(c) $E^{-}$, $C^{L}$, $a^{-}$]
Figure 5. Estimated trust state transition probabilities $\mathbb{P}(T_{t+1}|T_{t},E_{t+1},C_{t},a_{t})$ for low environment complexity $C_{t}=C^{L}$:
(a) $E^{+}$, $C^{H}$, $a^{+}$]
(b) $E^{-}$, $C^{H}$, $a^{+}$]
(c) $E^{+}$, $C^{H}$, $a^{-}$]
Figure 6. Estimated trust state transition probabilities $\mathbb{P}(T_{t+1}|T_{t},E_{t+1},C_{t},a_{t})$ for high environment complexity $C_{t}=C^{H}$:
4.1. Initial Trust and Observation Probabilities.
The estimated initial probability of trust being high is $\mathbb{P}(T^{H})=0.82$. This estimate suggests that at the beginning of the experiment, humans have a relatively high trust towards the robot.
@@ 4.3. Model Parameter Confidence
We now investigate uncertainty in the estimated parameters. Since the Baum-Welch algorithm provides a maximum a posteriori estimate, we rely on Laplace approximation (pml2Book, ) to estimate parameter uncertainty. To this end, we estimate the Hessian matrix $H$ of the log-likelihood $L$ and compute the uncertainty in parameter estimates as $e=\sqrt{\operatorname{diag}(-H)}$, where $\operatorname{diag}$ represents the diagonal elements of the matrix H-H. The uncertainties in the estimated IOHMM human trust model parameters are shown in Table 1.
@@
For parameters related to $C^{L}$, the standard errors are relatively low and accurate up to the first significant digit of the estimate. Specifically, this refers to the human action probability and $\mathbb{P}(T^{H}|T,E^{+},a^{+})$ parameters. However, the errors for other parameters are significantly higher, with the highest having an error of $\pm 0.56$. These high-uncertainty parameters are associated with scenarios involving failures and when the robot asks for assistance in low complexity which are attributed to the low frequency of these scenarios during the experiment.
The parameters conditioned on the state being in high trust $T^{H}$ have low standard errors, with the highest standard error being $\pm 0.07$. This suggests that these parameter estimates are accurate up to the first significant digit. In contrast, the standard errors for parameters conditioned on the state being in low trust $T^{L}$ are relatively high, ranging from $\pm 0.15$ to $\pm 0.3$. This high uncertainty can possibly be due to participants not getting into low trust state often, as compared to the high trust state.
@@ 4.4. Trust as a Hidden State
We now investigate if it is justified to refer to the hidden state in our model as “human trust”. During the experiment, after each subblock, participants completed a survey (lee1992trust, ) in which they responded, on an 11-point scale, to the following question “To what extent do you trust the automation in this scenario?”, shown in Appendix 8.2. This question was intended to capture participants’ level of trust based on their ongoing experience while performing the task. This survey is collected each time the robot finishes collecting objects from a shelf, before moving to another shelf.
Using the model parameters, we calculate the probability of being in a particular state at each time step via the belief update. Given the assumption that trust evolution follows a Markov process, the probability of trust being high is an aggregated representation of the trust state based on previous interactions. We compare the probability of trust being high at the end of each shelf-clearing event to the corresponding survey responses
The comparison of the probability of trust being high computed using the estimated model, and survey responses is shown in Fig. 8. It can be observed that the median probability of trust being high increases with trust scores collected through the survey. While the mid-level trust scores have higher variance, the higher trust estimates have lower variance in their trust response. Lower trust levels (0,1, and 2) did not receive any rating from participants. This can indicate that participants are more certain when their trust is high but uncertain about their trust level when it is low. The median of the probabilities of trust being high is used to fit a curve following a logistic function, resulting in
@@
where the function $P^{T}(r^{T})$ represents the probability of high trust as a function of the human trust response $r^{T}$. The curve is fitted using MATLAB’s fit function. The goodness of fit has been calculated using fit function. The Sum of Squares Error (SSE) is calculated to be equal to $0.0213$ and the Root Mean Squared Error (RMSE) is equal to $0.0653$. The plot of $P^{T}(r^{T})$ with the coefficients as compared with the median values of high trust estimate per survey response can be seen in Fig. 8. The results suggest that the hidden state in our model is isomorphic to self-reported trust scores, and the probability of trust being high is a sigmoidal function of the self-reported trust values. Thus, there exists a one-to-one mapping between trust estimates and self-reported survey responses, which suggests that the hidden state in our model is representative of human trust.
@@ 4.5. Trust and Human Reliance
It can be observed that the median survey responses can be approximated as a linear function of the trust survey response. The estimated function which maps trust response to reliance response can be approximated by the function
@@ 5.1. Computed Assistance-seeking Policy
The baseline policy that will be used is a trust-agnostic policy, which does not take the trust state into account. For the trust-agnostic policy, the only attribute of a trial is the environmental complexity. We used the data from the experiment in Section 4 to estimate the probability $\mathbb{P}(o_{t}=o^{-}|C_{t},a^{-})$, which is calculated by taking the empirical estimate in the dataset as
@@
$\mathbb{E}[R|a^{-},C]$ $=\mathcal{R}^{a^{+}}_{o^{+},E^{+}}\mathbb{P}(o_{t}=o^{+}|C_{t},a^{+})p^{\textup{suc}}_{C}+R^{a^{+}}_{o^{+},E^{-}}\mathbb{P}(o_{t}=o^{+}|C_{t},a^{+})(1-p^{\textup{suc}}_{C})+R^{a^{+}}_{o^{-}}\mathbb{P}(o_{t}=o^{-}|C_{t},a^{+}),$
@@
Since the complexity of the environment follows an i.i.d. random variable, for trust-agnostic policy, it suffices to simply calculate the expected reward per action for a given complexity to obtain the optimal action. With the data collected and our experiment design and parameters, the estimated rewards for attempting to collect autonomously are $\mathbb{E}[R|a^{+},C^{L}]=2.79$ and $\mathbb{E}[R|a^{+},C^{H}]=1.02$. Comparing the expected rewards for each action in the respective complexities, these calculations suggested that the optimal trust-agnostic policy is to never seek assistance irrespective of the environmental complexity, $a(t)=a^{+}$ for each $t$.
To highlight the importance of including task complexity in the assistance-seeking policy, we calculated a complexity-agnostic optimal policy. Specifically, we calculated an optimal policy that omits task complexity from its formulation. This is calculated by marginalizing over the complexity to obtain the state transition probability, which does not take complexity into account. The optimal action under this policy is to always collect autonomously, similar to the trust-agnostic policy.
@@ 5.2. Evaluation of the Proposed Policy
To evaluate the performance of the calculated optimal trust-aware assistance-seeking policy, a second set of experiments with human participants is conducted. A total of 20 participants were recruited and participated in the experiment, each completing two blocks of experiments in which the robot followed trust-aware and trust-agnostic policies, respectively. The participants recruited are different from those recruited in the first set of experiments. The order of the blocks was randomly assigned for each participant, with 10 participants starting with the trust-aware policy and the other 10 starting with the trust-agnostic policy. In each block, participants performed 20 low-complexity and 20 high-complexity trials. Due to the same reason as for the first experiment, data from 2 participants were was omitted, leaving us with 18 data sequences. The cumulative reward for all participants under both policies is illustrated in Fig. 11. The usage of trust-agnostic policy resulted in a median score of 69, while the trust-aware policy resulted in a median score of 84. The trust-aware policy collected more objects autonomously and prevented more human interruptions as compared with the trust-agnostic policy. A t-test with $5\%$ significance level shows that trust-aware policy outperforms trust-agnostic policy with a p-value of $p=0.0007$, and is calculated using MATLAB’s ttest2 function.
Figure 11. Cumulative reward statistics for the trust-agnostic and trust-aware policies. The trust-aware policy outperformed the trust-agnostic policy with a p-value of: 0.0007:
We ran additional experiments where participants’ compensation is dependent on the team’s performance. We informed participants that they will receive a base compensation of $\mathdollar 10$ and they can receive a bonus based on their aggregate performance/score. The average compensation received by the participants was $\mathdollar 13.18$. The results are shown in Fig. 13 in Appendix 8.3. The results are consistent with the above results, i.e., the trust-aware policy outperformed the trust-agnostic policy and the trust-aware policy received a higher trust score from participants.
5.3. Survey-Based Evaluation of the Trust-Aware Policy
@@
In total, participants completed this questionnaire three times, the first before the experiment, and the remaining two after completion with each policy. We dub these survey responses as “pre-experiment”, “post-trust-agnostic”, and “post-trust-aware” responses. The comparison of the survey scores is shown in Fig. 12. The responses to the 14 survey questions are aggregated, resulting in a trust score. This process involves first reverse coding the questions that describe negative qualities. The responses are then summed and divided by the total number of questions. More detailed instructions can be found in (schaefer2016measuring, ). The average scores on the collected trust survey are 65.44, 66.39, and 72.14 for pre-experiment, trust-agnostic, and trust-aware, respectively.
When comparing surveys collected after the experiment, the trust-aware policy (mean score of 72.14) yields a higher trust score relative to the trust-agnostic policy (mean score of 66.39) with a p-value of 0.0342, which is statistically significant. When comparing surveys collected before the experiment (mean score of 64.64) and post-trust-agnostic policy (mean score of 66.39), the p-value calculated is equal to 0.5973, which suggests that trust-agnostic policy has no statistically significant effect on trust. A comparison of the surveys collected before the experiment (mean score of 64.64) and after the experiment with the trust-aware policy (mean score of 72.14) suggests that the trust-aware policy improves human trust and the increase is statistically significant with a p-value of 0.0338. We have added a statistical summary of the collected data for this survey, which can be found in Appendix 8.1. In addition, to support the result, we have also collected trust and reliance survey responses after each subblock, and a statistical summary of the collected data can be seen in Appendix 8.2.
Figure 12. Comparison of the collected trust surveys before the experiment and after the experiment with the trust-aware and trust-agnostic policy. :